a 15 kg block slides on a horizontal surface with an initial speed of 12 m/s. the speed of 5.0m/s
is attained after travelling a distance of 25m.(a) what is the change in kinetic energy on the box (b) how much work was done by friction on the block ?
1
Expert's answer
2020-10-25T18:30:45-0400

a) The initial kinetic energy is:


"K_i = \dfrac{mv_i^2}{2}"

where "m=15kg" and "v_i = 12m\/s" is the initial speed of the block. Similarly, the final kinetic energy is:


"K_f = \dfrac{mv_f^2}{2}"

where "v_f = 5m\/s" is the final speed of the block.

The change is:


"\Delta K = K_f-K_i= \dfrac{mv_f^2}{2}- \dfrac{mv_i^2}{2} = \dfrac{m}{2}(v_f^2 - v_i^2)=\\
=\dfrac{15}{2}(5^2 - 12^2) = -892.5\space J"

b) According to the energy conservation law, the work done by friction is equal to the negative change of the kinetic energy:


"A = -\Delta K = 892.5\space J"

Answer. a) -892.5 J, b) 892.5 J.